The sum of the ages of a boy and his mother is 48. If the mother is three
more than twice the boy’s age, find the difference in the ages of the boy and
his mother.

Answer by fullofspark(1)

let assume the age of the boy be 'x' . so the age of mother would be '48-x' now according to the condition given in the question that
48-x = 2x+3
48-3=3x
x=15
you get the value of x that is 15
now the age of boy would be 15 and age of mother would be 33
so, difference would be 33-15= 18
18 would be your answer.